OpenStudy (anonymous):

Solve the equation below for x: log2x + log(2x-1) = log3x x = 5/4 x = 5/2 x = 0,5/4 There are no solutions.

OpenStudy (ranga):

Make use of the following identities: log(A) + log(B) = log(AB) If log(X) = log(Y) then X = Y

OpenStudy (ranga):

log2x + log(2x-1) = log(3x) log{ 2x(2x-1) } = log(3x) 2x(2x-1) = 3x 4x^2 - 2x - 3x = 0 4x^2 - 5x = 0 x(4x-5) = 0 x = 0 or x = 5/4 x = 0 is not allowed because log(0) is undefined. So x = 5/4